Water is flowing in a river at 2.5m/s. The river is 40 m wide with an
average depth of 6.0m. Compute the power available from the water
current
(PLEASE GIVE PROCESS)

Grade:9

1 Answers

Manas Satish Bedmutha
10 years ago
Consider time of 1 second. Volume of water flown = 6.0 x 40 x 2.5 = 600 m^3. Knowing density of water = 1000 kg/m3, Mass of water flowing = Vol. x Density = 600000 kg So, KE possessed by it = 0.5 x 600000 x 2.5 x 2.5 = 1875000 J. Hence power = Energy / Time = 1875000 / 1sec. = 1875000 W